OpenStudy (anonymous):

prove that arctan (x) + arctan (y) =(x+y)/(1-xy) (i have solved this part by using the tangent sum formula) the part i am stuck on is this use this relationship to show that if arctan (x) + arctan (y) + arctan (z) = pi/2 then xy + yz +zx =1

OpenStudy (anonymous):

after substitution and elimination i have got this equation' (x+y+z-xyz)/(1-xy-xz+yz) i don't know how to simplify it to reach xy+yz+zx

hartnn (hartnn):

i think there ia an error in denominator, u should get 1-xy-xz-yx then since arctan(..../....) is pi/2, u get (..../...) = infinity. that means denominator =0 1-xy-xz-yx=0 and u are done.... so just check your denominator

OpenStudy (anonymous):

ok thanks

hartnn (hartnn):

sorry, i meant 1-xy-zx-yz=0

OpenStudy (anonymous):

so do i move that to the right hand side

hartnn (hartnn):

1-xy-zx-yz=0 1=xy+zx+yz if u meant ^ then,yes.